Aphmau:

A football player kicks a 0.94 kg football with a force of 2.4 N. Calculate the acceleration of the football as the player kicks the ball. Round your answer to the nearest tenth of an m/s2. The acceleration of the football is m/s2.

Angle:

Hi @Aphmau ! Welcome to QuestionCove! Give me a minute to work out your problem :)

Aphmau:

ok

Angle:

I do believe there is a unit breakdown for Newtons \(1N = 1~kg~\times\large\frac{m}{s^2}\)

Angle:

we can replace the "1" idea with that one side of the equation should equal the other side it's like solving for missing variables: \(2.4 N = (0.94kg)\times(x ~\large\frac{m}{s^2})\) = 2.4 = 0.94 x solve for x

Angle:

@Aphmau let me know if you are still stuck or confused at all!

Aphmau:

ok thank you

YECR:

do u still need help?

Aphmau:

Yah